A father is 6 times older than his son in 9 years he will be 3 times older then his son how old is the son now?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

A father is 6 times older than his son in 9 years he will be 3 times older then his son how old is the son now?
------
Equations:
f = 6s
f+9 = 3(s+9)
---
Substitute for "f" and solve for "s":
6s+9 = 3s+27
3s = 18
s = 6 yrs old (the son's age now)
===================================
Cheers,
Stan H.
============

Answer by mananth(16946)

Son's age x
father's age 6x
9 years later
Son's age x+9
father's age 6x+9
father's age 3 times older
6x+9=3(x+9)
6x+9=3x+27
3x=27-9
3x=18
/3
x=6 Son's age
6*6=36 years father's age